# Lock-LLM: Prevent Unauthorized Knowledge Use from LLMs

### v1

@@ Abstract
The SmoothLLM defense provides a certification guarantee against jailbreaking attacks, but it relies on a strict ‘k-unstable’ assumption that rarely holds in practice. This strong assumption can limit the trustworthiness of the provided safety certificate. In this work, we address this limitation by introducing a more realistic probabilistic framework, ‘(k, $\varepsilon$)-unstable,’ to certify defenses against diverse jailbreaking attacks, from gradient-based (GCG) to semantic (PAIR). We derive a new, data-informed lower bound on SmoothLLM’s defense probability by incorporating empirical models of attack success, providing a more trustworthy and practical safety certificate. By introducing the notion of (k, $\varepsilon$)-unstable, our framework provides practitioners with actionable safety guarantees, enabling them to set certification thresholds that better reflect the real-world behavior of LLMs. Ultimately, this work contributes a practical and theoretically-grounded mechanism to make LLMs more resistant to the exploitation of their safety alignments, a critical challenge in secure AI deployment.
@@ 1 Introduction
Large Language Models (LLMs) face a critical vulnerability: "jailbreak" attacks that bypass safety protocols by manipulating input prompts to elicit objectionable responses zou2023universal; wei2023jailbreak. These attacks, ranging from gradient-based methods (GCG) to semantic approaches (PAIR) deng2023jailbreak, represent a fundamental challenge to LLM deployment.
SmoothLLM robey2023smoothllm provides the first formal certificate against jailbreak attacks by perturbing input prompts at the character level and aggregating responses. However, its certification relies on a strict "$k$-unstable" assumption: that adversarial prompts fail if any $k$ or more characters are altered. This deterministic assumption provides overly conservative guarantees that rarely hold in practice.
@@ 2 Related work
Early work demonstrated that aligned LLMs are vulnerable to carefully crafted prompts, with Wei et al. documenting instruction-following attack failures and Zou et al. introducing transferable methods like GCG that reliably coerce harmful outputs wei2023jailbroken; zou2023universal. The first formal certificate for jailbreak robustness is SmoothLLM robey2023smoothllm, which adapts randomized smoothing to the prompt space under a deterministic $k$-unstable assumption. Our work strengthens this theory by replacing the worst-case assumption with a probabilistic $(k,\varepsilon)$-unstable model and deriving tighter, data-driven bounds within the same smoothing framework.
@@ 3.1 Preliminaries: the SmoothLLM framework
Our work builds upon SmoothLLM robey2023smoothllm. Let $\text{LLM}(\cdot)$ denote an aligned large language model. An attacker crafts a prompt $P=[G;S]$ with goal string $G\in\mathcal{A}^{*}$ and adversarial suffix $S\in\mathcal{A}^{*}$ over alphabet $\mathcal{A}$ to:
@@ 3.2 Threat model
We assume a standard black-box threat model consistent with prior randomized defenses chiang2020certified; levine2020derandomized; cohen2019certified. The attacker cannot access random seeds for character perturbations or adapt their suffix in real-time based on the $N$ perturbed queries within a single SmoothLLM invocation (non-adaptive defense). This model aligns with Robey et al. robey2023smoothllm, who note that the non-differentiable nature of character-level perturbations poses significant challenges to adaptive attacks like GCG, despite surrogate-based adaptive attacks being theoretically possible.
@@ 3.3 Proposed probabilistic certification
Our proposed $(k,\varepsilon)$-unstable assumption is a direct generalization of the k-unstable assumption from robey2023smoothllm. By setting $\varepsilon=0$, we enforce that the probability of a jailbreak after $k$ or more character changes is zero. This recovers the original deterministic condition, where any such perturbation is guaranteed to neutralize the attack robey2023smoothllm. Our framework thus provides a more flexible model by accommodating the possibility of rare, successful perturbations.
@@ 3.5 Experimental motivation
Our experiments, conducted on Llama2 (7B) and Vicuna (7B) touvron2023llama; vicuna2023 with adversarial suffixes generated by the GCG attack and PAIR attack, reveal that this assumption is overly conservative. As shown in Figures 1 and 2, the ASR does not abruptly drop to zero but instead exhibits a clear exponential decay. The results clearly show a non-zero ASR that decays with k. This provides strong evidence against the strict k-unstable assumption and motivates the need for our proposed (k, $\varepsilon$)-unstable probabilistic framework, which is designed to model precisely this kind of empirically observed behavior. We provide extensive experimental analysis in Appendix E.
@@
All experiments were conducted on a single NVIDIA A100 GPU. We obtain 500 malicious prompts from zou2023universal and generate adversarial prompts using the Greedy Coordinate Gradient (GCG) attack and Prompt Automatic Iterative Refinement (PAIR) attack wei2023jailbroken; zou2023universal. Following the setup in the original SmoothLLM paper robey2023smoothllm, we evaluate two widely used open-source LLMs: Llama2 (7B) and Vicuna (7B) touvron2023llama; vicuna2023. Furthermore, we consider two perturbation strategies applied to the adversarial suffixes: RandomSwapPerturbation and RandomPatchPerturbation. We calculate Agresti-Coull confidence intervals since the attack success rate is extremely low.
@@ Appendix A Full proof of proposition 1
Note that the exponential decay model $\text{ASR}(i)\approx ae^{-bi}+c$ is theoretically motivated by the diminishing marginal impact of additional character perturbations. As more characters are modified, the probability of disrupting critical attack patterns increases exponentially, while the constant c captures residual success probability from attack robustness. This functional form has been validated across attack scenarios in adversarial ML literature [shafieinejadRobustnessBackdoorbasedWatermarking2019].
@@ Appendix D Combinatorics for patch perturbation
The number of starting positions that result in an overlap of size $i$ depends on whether the patch is fully contained within the goal or suffix, or if it straddles a boundary. Following the analysis framework of Robey et al. [robey2023smoothllm], we analyze the cases:
@@ Appendix E Additional experimental validation
This functional form is motivated by both empirical observations and prior work on adversarial suffix attacks. As shown in Figures 610, ASR decreases rapidly when only a few characters are perturbed and then levels off, indicating that adversarial suffixes often rely on a small subset of critical characters that drive the jailbreak logic. Perturbing these characters significantly disrupts the attack, while perturbing non-critical characters has little effect. This phenomenon is consistent with prior findings on GCG and PAIR jailbreaks, where specific trigger phrases or structural patterns are crucial to eliciting harmful outputs [zou2023universal, wei2023jailbroken].
@@
Our empirical results reveal a stark contrast in the robustness patterns of GCG and PAIR attacks, which can be attributed to their fundamentally different mechanisms. The GCG attack, which uses a gradient-based search to append an often nonsensical adversarial suffix [zou2023universal], demonstrates what can be termed syntactic fragility. Its success relies on a precise sequence of characters that exploits specific, brittle vulnerabilities in the model’s token processing. As shown in our experiments (e.g., Figures 2 and 6), this makes the attack highly susceptible to character-level perturbations; the ASR decays rapidly (a large fitted ‘b’ parameter) and approaches a near-zero floor (a small ‘c’). A few random character changes are often sufficient to break the exact “cheat code” the attack relies on.

### v2

@@ Abstract
The SmoothLLM defense provides a certification guarantee against jailbreaking attacks, but it relies on a strict ‘k-unstable’ assumption that rarely holds in practice. This strong assumption can limit the trustworthiness of the provided safety certificate. In this work, we address this limitation by introducing a more realistic probabilistic framework, ‘(k, $\varepsilon$)-unstable,’ to certify defenses against diverse jailbreaking attacks, from gradient-based (GCG) to semantic (PAIR). We derive a new, data-informed lower bound on SmoothLLM’s defense probability by incorporating empirical models of attack success, providing a more trustworthy and practical safety certificate. By introducing the notion of (k, $\varepsilon$)-unstable, our framework provides practitioners with actionable safety guarantees, enabling them to set certification thresholds that better reflect the real-world behavior of LLMs. Ultimately, this work contributes a practical and theoretically-grounded mechanism to make LLMs more resistant to the exploitation of their safety alignments, a critical challenge in secure AI deployment. Code is available at https://github.com/Adarsh321123/towards_realistic_guarantees.
@@ 1 Introduction
Large Language Models (LLMs) face a critical vulnerability: "jailbreak" attacks that bypass safety protocols by manipulating input prompts to elicit objectionable responses Zou et al. (2023); Wei et al. (2023a). These attacks, ranging from gradient-based methods (GCG) to semantic approaches (PAIR) Deng et al. (2024), represent a fundamental challenge to LLM deployment.
SmoothLLM Robey et al. (2023) provides the first formal certificate against jailbreak attacks by perturbing input prompts at the character level and aggregating responses. However, its certification relies on a strict "$k$-unstable" assumption: that adversarial prompts fail if any $k$ or more characters are altered. This deterministic assumption provides overly conservative guarantees that rarely hold in practice.
@@ 2 Related work
Early work demonstrated that aligned LLMs are vulnerable to carefully crafted prompts, with Wei et al. documenting instruction-following attack failures and Zou et al. introducing transferable methods like GCG that reliably coerce harmful outputs Wei et al. (2023b); Zou et al. (2023). The first formal certificate for jailbreak robustness is SmoothLLM Robey et al. (2023), which adapts randomized smoothing to the prompt space under a deterministic $k$-unstable assumption. Our work strengthens this theory by replacing the worst-case assumption with a probabilistic $(k,\varepsilon)$-unstable model and deriving tighter, data-driven bounds within the same smoothing framework.
@@ 3.1 Preliminaries: the SmoothLLM framework
Our work builds upon SmoothLLM Robey et al. (2023). Let $\text{LLM}(\cdot)$ denote an aligned large language model. An attacker crafts a prompt $P=[G;S]$ with goal string $G\in\mathcal{A}^{*}$ and adversarial suffix $S\in\mathcal{A}^{*}$ over alphabet $\mathcal{A}$ to:
@@ 3.2 Threat model
We assume a standard black-box threat model consistent with prior randomized defenses Chiang et al. (2020); Levine and Feizi (2020); Cohen et al. (2019). The attacker cannot access random seeds for character perturbations or adapt their suffix in real-time based on the $N$ perturbed queries within a single SmoothLLM invocation (non-adaptive defense). This model aligns with Robey et al. Robey et al. (2023), who note that the non-differentiable nature of character-level perturbations poses significant challenges to adaptive attacks like GCG, despite surrogate-based adaptive attacks being theoretically possible.
@@ 3.3 Proposed probabilistic certification
Our proposed $(k,\varepsilon)$-unstable assumption is a direct generalization of the k-unstable assumption from Robey et al. (2023). By setting $\varepsilon=0$, we enforce that the probability of a jailbreak after $k$ or more character changes is zero. This recovers the original deterministic condition, where any such perturbation is guaranteed to neutralize the attack Robey et al. (2023). Our framework thus provides a more flexible model by accommodating the possibility of rare, successful perturbations.
@@ 3.5 Experimental motivation
Our experiments, conducted on Llama2 (7B) and Vicuna (7B) Touvron et al. (2023); Chiang et al. (2023) with adversarial suffixes generated by the GCG attack and PAIR attack, reveal that this assumption is overly conservative. As shown in Figures 1 and 2, the ASR does not abruptly drop to zero but instead exhibits a clear exponential decay. The results clearly show a non-zero ASR that decays with k. This provides strong evidence against the strict k-unstable assumption and motivates the need for our proposed (k, $\varepsilon$)-unstable probabilistic framework, which is designed to model precisely this kind of empirically observed behavior. We provide extensive experimental analysis in Appendix E.
@@
All experiments were conducted on a single NVIDIA A100 GPU. We obtain 500 malicious prompts from Zou et al. (2023) and generate adversarial prompts using the Greedy Coordinate Gradient (GCG) attack and Prompt Automatic Iterative Refinement (PAIR) attack Wei et al. (2023b); Zou et al. (2023). Following the setup in the original SmoothLLM paper Robey et al. (2023), we evaluate two widely used open-source LLMs: Llama2 (7B) and Vicuna (7B) Touvron et al. (2023); Chiang et al. (2023). Furthermore, we consider two perturbation strategies applied to the adversarial suffixes: RandomSwapPerturbation and RandomPatchPerturbation. We calculate Agresti-Coull confidence intervals since the attack success rate is extremely low.
@@ Appendix A Full proof of proposition 1
Note that the exponential decay model $\text{ASR}(i)\approx ae^{-bi}+c$ is theoretically motivated by the diminishing marginal impact of additional character perturbations. As more characters are modified, the probability of disrupting critical attack patterns increases exponentially, while the constant c captures residual success probability from attack robustness. This functional form has been validated across attack scenarios in adversarial ML literature [7].
@@ Appendix D Combinatorics for patch perturbation
The number of starting positions that result in an overlap of size $i$ depends on whether the patch is fully contained within the goal or suffix, or if it straddles a boundary. Following the analysis framework of Robey et al. [6], we analyze the cases:
@@ Appendix E Additional experimental validation
This functional form is motivated by both empirical observations and prior work on adversarial suffix attacks. As shown in Figures 610, ASR decreases rapidly when only a few characters are perturbed and then levels off, indicating that adversarial suffixes often rely on a small subset of critical characters that drive the jailbreak logic. Perturbing these characters significantly disrupts the attack, while perturbing non-critical characters has little effect. This phenomenon is consistent with prior findings on GCG and PAIR jailbreaks, where specific trigger phrases or structural patterns are crucial to eliciting harmful outputs [11, 10].
@@
Our empirical results reveal a stark contrast in the robustness patterns of GCG and PAIR attacks, which can be attributed to their fundamentally different mechanisms. The GCG attack, which uses a gradient-based search to append an often nonsensical adversarial suffix [11], demonstrates what can be termed syntactic fragility. Its success relies on a precise sequence of characters that exploits specific, brittle vulnerabilities in the model’s token processing. As shown in our experiments (e.g., Figures 2 and 6), this makes the attack highly susceptible to character-level perturbations; the ASR decays rapidly (a large fitted ‘b’ parameter) and approaches a near-zero floor (a small ‘c’). A few random character changes are often sufficient to break the exact “cheat code” the attack relies on.